OpenStudy (anonymous):

Don't think I am doing this right. 4(w+7)-6=11(w+2)

OpenStudy (anonymous):

www.wolframalpha.com <-------will help you will all your problems

OpenStudy (anonymous):

its a math site super usefull tool.

OpenStudy (anonymous):

asumeing you are being told to solve for W 4(w+7)-6=11(w+2) W=0

OpenStudy (anonymous):

plug 0 in and add up both sides and you know i am correct if it comes out where one side is equal tot he other.

OpenStudy (anonymous):

to the other*

OpenStudy (anonymous):

it came out to 1=2

OpenStudy (anonymous):

hmm recheck im almost positive its 0

OpenStudy (anonymous):

w=0 :)

OpenStudy (anonymous):

oh never mind your right w=0

OpenStudy (phi):

4(w+7)-6=11(w+2) distribute the 4 which means multiply 4 times each term inside the parens 4w + 4*7 - 6 = 11(w + 2) distribute the 11 on the right side 4w + 28 - 6 = 11w + 22 simplify 28-6 to 22 4w + 22 = 11w + 22 add -22 to both sides 4w + 22 -22 = 11w + 22 -22 22 -22 is zero, so this is 4w = 11w add -4w to both sides 4w - 4w = 11w - 4w 0 = 7w divide both sides by 7 0/7 = 7w/7 or 0 = w

OpenStudy (anonymous):

THANK YOU!!!!

OpenStudy (anonymous):

yay! np

OpenStudy (anonymous):

I also have this one but idk if i did it right either 5(3x+12)=2(8x-16) i did 5(3x)+5(12)=2(8x)-2(-16) so 15x+60=16x+32 -32 -32 15x+28=16x -15 -15 28=1x 28=x

OpenStudy (anonymous):

i also did it another way and got 92

OpenStudy (anonymous):

plug in 28 to both sides and check if both numbers on each side of the = sign are same then you are correct.

OpenStudy (anonymous):

yeah when i do that it never works out right idk why iv done this before but its just not working out so well for me today

OpenStudy (phi):

you have to be careful of the minus signs 5(3x+12)=2(8x-16) i did 5(3x)+5(12)=2(8x)-2(-16) notice on the right side of the equation that 2(8x-16) can be written as 2(8x + -16) It is less confusing if we DON'T USE SUBTRACTION. instead add negative numbers. now you distribute the 2: 2*8x + 2*-16 you get 16x -32 which is different from what you got.

OpenStudy (phi):

so you now have 5(3x)+5(12) = 2(8x)+ 2(-16) can you finish ?

OpenStudy (anonymous):

oh okay I get it thank you!